# Development & Characterization of Electrodes for large-scale Xenon Time Projection Chambers

### v1

@@ 2.3 Wire Installation Procedure
In previous experiments, the common strategy of mounting several hundreds of parallel wires onto an electrode frame was a challenging and iterative process [lz_electrodes, mannino_thesis]. For each newly installed wire, the electrode frame deforms simultaneously, altering the tension of those wires already mounted. For an even larger electrode, a new installation method is required to reduce the number of individual steps, the time needed, and potential inaccuracies during the process. In this work, we developed a new set of tools and procedures for a robust parallel-wire electrode installation, as shown in Figure 2. Essentially, we deformed the electrode frame to the final equilibrium shape before mounting the wires. Thus, the frame held its shape throughout the wire installation. Afterwards, the assembled electrode was dismounted from the station.
@@
During wire installation, the shape of the electrode frame underwent limited change due to the stiffness of the tensioning rods, amounting to $0.35\pm 0.24$ $\text{\,}\mathrm{m}\mathrm{m}$. The wires were installed with their target tension values as indicated in Figure 2. With custom-made tools, we tensioned the wire on both ends through the pin hole and fixed the wire using copper pins. All individual wire tensions were measured by a laser system pioneered in [Prall:2009ep_laser_system], updated and fabricated by [Jens2023].
@@ 3 Hexagonal Mesh Electrodes
The meshes were produced with openings measuring $7.5\text{\,}\mathrm{m}\mathrm{m}$ and leg thickness of $0.3\text{\,}\mathrm{m}\mathrm{m}$. The choice of leg thickness was driven by exploratory electropolishing which was conducted with test meshes with leg cross-section of $0.2\text{\,}\mathrm{m}\mathrm{m}$ at OTG Oberflächentechnik Gronau [OTG]. After electropolishing these test meshes exhibited loss of material at the outer radii which was significant enough to cause breaks in individual hex legs. Therefore, to minimize the number of potentially cut hex legs, we opted for thicker mesh legs of $0.3\text{\,}\mathrm{m}\mathrm{m}$. Additionally, attempts were made to investigate the potential use of water-jet cutting as an alternative to photochemical etching. However, preliminary results suggested that photochemical etching was superior in terms of production time, material use and the resulting mesh quality.
Each of the half-meshes has an outer and an inner ring for stretching and fixation of the mesh on the electrode frame, respectively. The rings are interconnected via narrow strips, which are intended to be mechanically severed after the mesh is stretched and fixed to the electrode frame. Additionally, omega-shaped relief sections are located at regularly spaced gaps along both the inner and outer rings, facilitating greater flexibility of both rings. Four of the produced half-meshes were laser-welded along a dedicated $2\text{\,}\mathrm{m}\mathrm{m}$ joining edge on each half, resulting in a central welded strip of $4\text{\,}\mathrm{m}\mathrm{m}$ across the final two whole meshes. The strength of the connection produced by the laser welding was assessed in a series of tests that will be further elaborated on in subsection 3.4. A ring-shaped SS electrode frame with an outer diameter of $1395\text{\,}\mathrm{m}\mathrm{m}$ and thickness of $20\text{\,}\mathrm{m}\mathrm{m}$ was produced and electropolished by MPS [MPS]. To host the electrode mesh, the ring has a semi-circular cross-section and features 96 threaded M4-sized holes through which the mesh could be fastened to the frame.
@@ 3.1 Mechanical, Optical & Electrostatic Simulations
Given the high field between the anode and the gate in the gaseous phase, inhomogeneity in the electrode’s geometry largely affects the S2 signal. Therefore, the design with the central welded strip is not the most suitable candidate for the anode and the gate electrodes. The subsequent discussion then focuses solely on the cathode electrode. To study the effect of the $4\text{\,}\mathrm{m}\mathrm{m}$ welded strip on the field homogeneity, we performed both a full 3D and a local 3D electrostatic field simulation. We used the KEMField package within the open-source software Kassiopeia [Kassiopeia] for the full 3D simulation. The package uses the boundary element method for the surface charge calculation and the fast multipole method to perform the field calculation. Since the hexagonal geometry is computationally expensive, we used wire grids together with a $4\text{\,}\mathrm{m}\mathrm{m}$ wide plate on top of the grid to represent the electrode.
@@
Subsequently, a local 3D simulation with finer discretization and accurate geometry was performed using the Electrostatics module of COMSOL Multiphysics® software [COMSOL]. The simulation domain is shown in Figure 7 (left). The hexagonal mesh geometry was imported from the CAD drawing. The top and bottom boundary conditions were retrieved from the aforementioned full 3D field simulation. Zero-charged boundary conditions were imposed on the four sides, meaning the simulation domain repeated infinitely.
@@
To evaluate the result in both simulations, we consider an array of lines perpendicular to the welded strip at $y=0$ and $x\in[-40,40]$ $\text{\,}\mathrm{m}\mathrm{m}$, parallel in z-direction. The normalized standard deviation of the field magnitude along these lines was used as the metric for field homogeneity. In comparison, the same calculation was performed for parallel wires without the central welded strip. The result is shown in Figure 7. Compared to the parallel wire case, the case with the welded strip has a larger deviation in the field as expected. The closer it is to the cathode electrode, the larger the field inhomogeneity. However, the break even point is around 7.5 cm above the cathode, and is comparable to the previous XENONnT fiducial volume bottom boundary [nT_analysis_SR0:2025].
@@
Figure 7: Left: close-up of the result of the local 3D simulation with COMSOL®. The welded strip is along the y-direction. The $E_{norm}$ field on the x-z plane at $y=0$ is also shown by the colour map. Right: the standard deviation of $E_{norm}$ for $x\in[-40,40]$ $\text{\,}\,\mathrm{m}\mathrm{m}$ normalized by the mean field at the same z-distance from the cathode. The field inhomogeneity due to the welded strip becomes negligible at 7.5 cm above the cathode. Results from both the local 3D (solid line) and global 3D (dashed line) simulations from COMSOL® and KEMField are shown. The color indicates the electrode configuration, with a hexagonal mesh cathode in blue and a parallel wire cathode in orange. The fiducial volume cut used in a previous XENONnT analysis [nT_analysis_SR0:2025] is also indicated by the dashed-dotted line.
@@
The electrostatic force on the cathode was derived using a similar COMSOL local 3D simulation with an infinitely large cathode, but without the welded strip in place. The resulting load was $1.5\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}^{2}$. Additional loads, namely tensioning stress, gravitational loads, and thermal shrinkage, were also considered for the mechanical simulation. Similar to subsection 2.1, the ANSYS Workbench 2021 was used to determine the stress and the safety factor of the hexagonal mesh under loads. To reduce the computational time, a 4-fold mirror symmetry and a scale factor of 2 were applied to the model. With a simpler model, we verified that the scale factor did not change the result as long as the mass of the electrode remained unchanged.
@@
From the result, the maximum stress is at $167\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$ with a $25\text{\,}\mathrm{K}$ temperature difference between the mesh and the frame. This corresponds to a safety factor of 1.4, assuming the YTS for the material to be at $240\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$. The expected sagging of the mesh during operation is well below the maximally allowed sagging of 2 mm as shown in subsection 3.2.
@@
Comparing the mentioned hexagonal mesh of the leg width of $0.3\text{\,}\mathrm{m}\mathrm{m}$, an opening of $7.5\text{\,}\mathrm{m}\mathrm{m}$ and a $4\text{\,}\mathrm{m}\mathrm{m}$-welded strip at the center, to a parallel-wire electrode with $0.3\text{\,}\mathrm{m}\mathrm{m}$ diameter wires with the pitch of $7.5\text{\,}\mathrm{m}\mathrm{m}$, the transparency is lowered by 4%. This can be used as a proxy for the first-order effect on the Light Collection Efficiency (LCE): the probability of detecting a photon produced in the TPC.
@@
A second-order effect is the impact of the shadow cast by the electrode onto the PMT array. To study this electrode shadow, we set up a Monte-Carlo simulation of a complete XLZD-scale TPC using GEANT4 [geant4]. We implemented four different electrode designs: An opaque disk with transparency equal to the transparent area of a typical electrode, an electrode made of parallel wires, one made of parallel and perpendicular wires forming a square pattern, and an electrode with a hexagonal mesh pattern.
@@ 3.2 Tensioning & Assembly Mechanism
After the stretching process was complete, the mesh was fixed to the electrode frame using dedicated 24 clamping pieces, and vented M4 Ag-coated screws. Then, all the ancillary segments were removed and the stretching ring was cut off by severing the connections between the stretching ring and mesh’s inner ring. To conceal the cut area and the clamps, additional 24 cover pieces were placed on top of each clamp and fixed with vented M2.5 Au-coated screws. The assembly process was first tested at KIT in non-cleanroom conditions, and then was carried out inside a cleanroom at LNGS. The sagging of the mesh throughout the stretching process was verified by measuring the distance between the mesh’s surface and a reference surface at the center of the SS mounting plate. The sagging of the mesh was controlled at every stage of the installation until mounting the electrode in the detector. The measured sagging of 0.6 mm is expected to lead to a maximum sagging of 1.1 mm during operation, well below the requirement of 2 mm.
@@ 3.3 ML-based Defect Detection
During production or handling, a mesh can be damaged, potentially deteriorating the mechanical stability of the mesh, or locally altering the produced electric field eventually leading to electron emission. To be able to assess the impact of different features, such as the ones seen in Figure 10, they must first be identified. A typical mesh for a large-scale electrode can consist of hundreds of thousands of individual legs, making a full inspection by eye unfeasible. Reconstruction based anomaly detection using Machine Learning (ML) models can be used to find physical defects in images of workpieces [yolov5-defect-detection, cascading-ae-defect, granite_optical].
@@
In a VAE, the latent variable $z$ is replaced by $m$ uncorrelated Normal Distributions, parametrized by a mean value $\vec{\mu}\in\mathbb{R}^{m}$ and a standard deviation $\vec{\sigma}\in\mathbb{R}_{+}^{m}$, which are learned by the encoder. The decoder then samples from the input-dependent multivariate Normal Distribution $z\sim\mathcal{N}(\vec{\mu},\vec{\sigma}|x)$. The VAE in this work was trained on the Evidence Lower Bound loss function, as is usual for such models [vae-introduction]. The performance of the VAE was evaluated on 4000 clean legs and 4000 digitally augmented images of defect legs. Figure 12 shows the Receiver-Operating-Characteristic (ROC) of the VAE model, where the positive case corresponds to an image being identified as showing a feature. The Area Under the Curve of the VAEs ROC is $\text{AUC}=0.98$, which is comparable to other defect detection models [mem-ae].
@@ 4 High-Voltage Testing & Performance
Features, such as the ones showcased in subsection 3.3, could lead to localized field enhancement, which in turn could result in detrimental effects such as field emission and even breakdown [adam_bailey_thesis_em, pixie]. To assess the performance of an electrode, one cannot rely solely on electrostatic simulations, as the field around each feature is highly dependent on its geometry. Hence, to study such effects and to determine the overall performance of the electrode, we have performed a series of HV tests in a gaseous argon (GAr) environment. In those tests, we aimed to observe light associated with effects such as field emission or breakdown events.
@@ 4.1 High-Voltage Test Setup
A test setup, inspired by the one described in [daniel_wenz_thesis, Mitra2023], was built inside the Faraday cage. As illustrated in Figure 13, at the heart of the setup lies a square PVC box measuring $1.6\text{\,}\mathrm{m}$ in length and $20\text{\,}\mathrm{c}\mathrm{m}$ in depth. The box was elevated above the floor with the use of a dedicated four-legged frame with wheels. To maintain a GAr atmosphere and protect the electrode under test from environmental contaminants the box was instrumented with a 1.1 $1\text{\,}\mathrm{c}\mathrm{m}$ thick lid made of Makroclear® polycarbonate sheet [makroclear], with strips of Armaflex® tape for sealing. The use of the polycarbonate lid transmits the light in the visible and in the near-infrared range [makroclear_transmittance, polycarbonate_transmittance:2018], allowing us to observe light, which could be associated with effects such as neutral bremsstrahlung electroluminescence and the avalanche scintillation of GAr in the near-infrared range [ELreview2020, gAr_spectrum:2022]. The box was fitted with a GAr tube inlet, a HV connection to the electrode and a ground connection. The flow of GAr into the box was controlled via a flow meter connected to the bottles containing grade 5.0 argon. Additionally, an Infactory NC-7004-675 humidity and temperature sensor [infactory] was placed at the farthest corner from the GAr inlet, as these parameters can affect the breakdown voltage. Since the PVC box was not leak-tight, the pressure inside the box was assumed to be close to the atmospheric pressure, and thus not monitored.
@@
A $1.2\text{\,}\mathrm{m}$ in diameter and $5\text{\,}\mathrm{m}\mathrm{m}$ thick matt aluminum plate with a rounded edge was placed at the bottom of the box on top of a grounded SS plate, forming a ground plate in the setup. A number of PVC spacers was used to further support the aluminum plate on its sides, resulting in a total estimated elevation of $1\text{\,}\mathrm{c}\mathrm{m}$ of the plate above the box’s bottom. The maximum measured gap distance in between the electrode under test and the ground plate was $4.97\text{\,}\mathrm{c}\mathrm{m}$. This measured gap distance was used to define the so-called bulk field in subsequent analysis. However, it should be noted that throughout the HV tests the precise distance and alignment between the ground plate and the tested electrode were not controlled. The unevenness of the ground plate, defined as the vertical distance between its highest and lowest points, was measured to be $6.4\text{\,}\mathrm{m}\mathrm{m}$, amounting to about 13% of the gap distance.
The electrode mesh was mounted on the electrode frame in non-cleanroom environment conditions, with the use of the procedure described in subsection 3.2. However, as depicted in Figure 13 (top), for the purpose of the HV tests the stretching ring was not removed and the SS covers were not installed. The electrode was positioned in the setup with the side on which the mesh was mounted facing down towards the ground plate. To facilitate the HV connection, the electrode was connected to a connection in the box’s wall, which subsequently was connected to the toroid. For most of the tests, $20\text{\,}\mathrm{M\SIUnitSymbolOhm}$ resistors were added on the HV line leading from the test box to the toroid, limiting the current into the system and preventing potential electrode aging effects as a result of discharges [electrode_againg_1973]. In the final test configuration, the electrode frame was suspended from the box’s lid using PEEK [insulator:2018] pillars wrapped in Kapton tape, as illustrated in Figure 13. Prior to that, several insulators with a varying geometry, produced from PVC, POM, and 3D-printed ASA plastic were tested as separators between the electrode frame and the ground plate. However, their use resulted in frequent breakdowns at the insulators, indicating that these insulators enhanced locally the field along the edge of the ground plate.

### v2

@@ 2.3 Wire Installation Procedure
In previous experiments, the common strategy of mounting several hundreds of parallel wires onto an electrode frame was a challenging and iterative process [56, 57]. For each newly installed wire, the electrode frame deforms simultaneously, altering the tension of those wires already mounted. For an even larger electrode, a new installation method is required to reduce the number of individual steps, the time needed, and potential inaccuracies during the process. In this work, we developed a new set of tools and procedures for a robust parallel-wire electrode installation, as shown in Figure 2. Essentially, we deformed the electrode frame to the final equilibrium shape before mounting the wires. Thus, the frame held its shape throughout the wire installation. Afterwards, the assembled electrode was dismounted from the station.
@@
During wire installation, the shape of the electrode frame underwent limited change due to the stiffness of the tensioning rods, amounting to $0.35\pm 0.24$ $\text{\,}\mathrm{m}\mathrm{m}$. The wires were installed with their target tension values as indicated in Figure 2. With custom-made tools, we tensioned the wire on both ends through the pin hole and fixed the wire using copper pins. All individual wire tensions were measured by a laser system pioneered in [70], updated and fabricated by [74].
@@ 3 Hexagonal Mesh Electrodes
The meshes were produced with openings measuring $7.5\text{\,}\mathrm{m}\mathrm{m}$ and leg thickness of $0.3\text{\,}\mathrm{m}\mathrm{m}$. The choice of leg thickness was driven by exploratory electropolishing which was conducted with test meshes with leg cross-section of $0.2\text{\,}\mathrm{m}\mathrm{m}$ at OTG Oberflächentechnik Gronau [68]. After electropolishing these test meshes exhibited loss of material at the outer radii which was significant enough to cause breaks in individual hex legs. Therefore, to minimize the number of potentially cut hex legs, we opted for thicker mesh legs of $0.3\text{\,}\mathrm{m}\mathrm{m}$. Additionally, attempts were made to investigate the potential use of water-jet cutting as an alternative to photochemical etching. However, preliminary results suggested that photochemical etching was superior in terms of production time, material use and the resulting mesh quality.
Each of the half-meshes has an outer and an inner ring for stretching and fixation of the mesh on the electrode frame, respectively. The rings are interconnected via narrow strips, which are intended to be mechanically severed after the mesh is stretched and fixed to the electrode frame. Additionally, omega-shaped relief sections are located at regularly spaced gaps along both the inner and outer rings, facilitating greater flexibility of both rings. Four of the produced half-meshes were laser-welded along a dedicated $2\text{\,}\mathrm{m}\mathrm{m}$ joining edge on each half, resulting in a central welded strip of $4\text{\,}\mathrm{m}\mathrm{m}$ across the final two whole meshes. The strength of the connection produced by the laser welding was assessed in a series of tests that will be further elaborated on in subsection 3.4. A ring-shaped SS electrode frame with an outer diameter of $1395\text{\,}\mathrm{m}\mathrm{m}$ and thickness of $20\text{\,}\mathrm{m}\mathrm{m}$ was produced and electropolished by MPS [63]. To host the electrode mesh, the ring has a semi-circular cross-section and features 96 threaded M4-sized holes through which the mesh could be fastened to the frame.
@@ 3.1 Mechanical, Optical & Electrostatic Simulations
Given the high field between the anode and the gate in the gaseous phase, inhomogeneity in the electrode’s geometry largely affects the S2 signal. Therefore, the design with the central welded strip is not the most suitable candidate for the anode and the gate electrodes. The subsequent discussion then focuses solely on the cathode electrode with its requirements as stated in Table 4. To study the effect of the $4\text{\,}\mathrm{m}\mathrm{m}$ welded strip on the field homogeneity, we performed both a full 3D and a local 3D electrostatic field simulation. We used the KEMField package within the open-source software Kassiopeia [43] for the full 3D simulation. The package uses the boundary element method for the surface charge calculation and the fast multipole method to perform the field calculation. Since the hexagonal geometry is computationally expensive, we used wire grids together with a $4\text{\,}\mathrm{m}\mathrm{m}$ wide plate on top of the grid to represent the electrode.
@@
Subsequently, a local 3D simulation with finer discretization and accurate geometry was performed using the Electrostatics module of COMSOL Multiphysics® software [36]. The simulation domain is shown in Figure 7 (left). The hexagonal mesh geometry was imported from the CAD drawing. The top and bottom boundary conditions were retrieved from the aforementioned full 3D field simulation. Zero-charged boundary conditions were imposed on the four sides, meaning the simulation domain repeated infinitely.
Table 4: Requirements for the mesh electrode as a cathode. The maximum sagging of the mesh under electrostatic and gravitational force ($h_{\text{max}}$) was estimated based on the electrode configuration in the XENONnT detector. The yield tensile strength for the mesh ($\text{YTS}_{0.2\%}$) was taken as the typical value of SS304 from [38]. The thermal stress ($\sigma_{s}$) is defined as in Table 1. The stress due to the tension applied to the mesh is defined as $\sigma_{T}$.
Quantity Symbol Requirement
Sagging limit $h_{\text{max}}$ $2.0\text{\,}\mathrm{m}\mathrm{m}$
Safety factor $\text{YTS}_{0.2\%}\,/\,(\sigma_{T}+\sigma_{s})$ $\gtrsim 1.5$
Transparency $\gtrsim 90\%$
@@
To evaluate the result in both simulations, we consider an array of lines perpendicular to the welded strip at $y=0$ and $x\in[-40,40]$ $\text{\,}\mathrm{m}\mathrm{m}$, parallel in z-direction. The normalized standard deviation of the field magnitude along these lines was used as the metric for field homogeneity. In comparison, the same calculation was performed for parallel wires without the central welded strip. The result is shown in Figure 7. Compared to the parallel wire case, the case with the welded strip has a larger deviation in the field as expected. The closer it is to the cathode electrode, the larger the field inhomogeneity. However, the break even point is around 7.5 cm above the cathode, and is comparable to the previous XENONnT fiducial volume bottom boundary [26].
@@
Figure 7: Left: close-up of the result of the local 3D simulation with COMSOL®. The welded strip is along the y-direction. The $E_{norm}$ field on the x-z plane at $y=0$ is also shown by the colour map. Right: the standard deviation of $E_{norm}$ for $x\in[-40,40]$ $\text{\,}\,\mathrm{m}\mathrm{m}$ normalized by the mean field at the same z-distance from the cathode. The field inhomogeneity due to the welded strip becomes negligible at 7.5 cm above the cathode. Results from both the local 3D (solid line) and global 3D (dashed line) simulations from COMSOL® and KEMField are shown. The color indicates the electrode configuration, with a hexagonal mesh cathode in blue and a parallel wire cathode in orange. The fiducial volume cut used in a previous XENONnT analysis [26] is also indicated by the dashed-dotted line.
@@
The electrostatic force on the cathode was derived using a similar COMSOL local 3D simulation with an infinitely large cathode, but without the welded strip in place. The resulting load was $1.5\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}^{2}$. Additional loads, namely tensioning stress, gravitational loads, and thermal shrinkage, were also considered for the mechanical simulation. Similar to subsection 2.1, the ANSYS Workbench 2021 was used to determine the stress and the safety factor of the hexagonal mesh under loads. To reduce the computational time, a 2-fold mirror symmetry and a scale factor of 2 were applied to the model. With a simpler model, we verified that the scale factor did not change the result as long as the mass of the electrode remained unchanged.
@@
From the result, the maximum stress ($\sigma_{T}+\sigma_{s}$) is at $167\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$ with a $25\text{\,}\mathrm{K}$ temperature difference between the mesh and the frame. This corresponds to a safety factor of 1.4, assuming the YTS for the material to be at $240\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$ [38]. The expected sagging of the mesh during operation is well below the maximally allowed sagging of 2 mm as shown in subsection 3.2.
@@
Comparing the mentioned hexagonal mesh of the leg width of $0.3\text{\,}\mathrm{m}\mathrm{m}$, an opening of $7.5\text{\,}\mathrm{m}\mathrm{m}$ and a $4\text{\,}\mathrm{m}\mathrm{m}$-welded strip at the center, to a parallel-wire electrode with $0.3\text{\,}\mathrm{m}\mathrm{m}$ diameter wires with the pitch of $7.5\text{\,}\mathrm{m}\mathrm{m}$, the transparency is lowered by 4% reaching $91.4\%$. This can be used as a proxy for the first-order effect on the Light Collection Efficiency (LCE): the probability of detecting a photon produced in the TPC.
@@
A second-order effect is the impact of the shadow cast by the electrode onto the PMT array. To study this electrode shadow, we set up a Monte-Carlo simulation of a complete XLZD-scale TPC using GEANT4 [10]. We implemented four different electrode designs: An opaque disk with transparency equal to the transparent area of a typical electrode, an electrode made of parallel wires, one made of parallel and perpendicular wires forming a square pattern, and an electrode with a hexagonal mesh pattern.
@@ 3.2 Tensioning & Assembly Mechanism
After the stretching process was complete, the mesh was fixed to the electrode frame using dedicated 24 clamping pieces, and vented M4 Ag-coated screws. Then, all the ancillary segments were removed and the stretching ring was cut off by severing the connections between the stretching ring and mesh’s inner ring. To conceal the cut area and the clamps, additional 24 cover pieces were placed on top of each clamp and fixed with vented M2.5 Au-coated screws. The assembly process was first tested at KIT in non-cleanroom conditions, and then was carried out inside a cleanroom at LNGS. The sagging of the mesh throughout the stretching process was verified by measuring the distance between the mesh’s surface and a reference surface at the center of the SS mounting plate. The sagging of the mesh was controlled at every stage of the installation until mounting the electrode in the detector. The measured sagging of 0.6 mm is expected to lead to a maximum sagging $h_{\text{max}}$ of 1.1 mm during operation, well below the requirement of 2 mm, as given in Table 4.
@@ 3.3 ML-based Defect Detection
During production or handling, a mesh can be damaged, potentially deteriorating the mechanical stability of the mesh, or locally altering the produced electric field eventually leading to electron emission. To be able to assess the impact of different features, such as the ones seen in Figure 10, they must first be identified. A typical mesh for a large-scale electrode can consist of hundreds of thousands of individual legs, making a full inspection by eye unfeasible. Reconstruction based anomaly detection using Machine Learning (ML) models can be used to find physical defects in images of workpieces [35, 80, 39].
@@
In a VAE, the latent variable $z$ is replaced by $m$ uncorrelated Normal Distributions, parametrized by a mean value $\vec{\mu}\in\mathbb{R}^{m}$ and a standard deviation $\vec{\sigma}\in\mathbb{R}_{+}^{m}$, which are learned by the encoder. The decoder then samples from the input-dependent multivariate Normal Distribution $z\sim\mathcal{N}(\vec{\mu},\vec{\sigma}|x)$. The VAE in this work was trained on the Evidence Lower Bound loss function, as is usual for such models [52]. The performance of the VAE was evaluated on 4000 clean legs and 4000 digitally augmented images of defect legs. Figure 12 shows the Receiver-Operating-Characteristic (ROC) of the VAE model, where the positive case corresponds to an image being identified as showing a feature. The Area Under the Curve of the VAEs ROC is $\text{AUC}=0.98$, which is comparable to other defect detection models [45].
@@ 4 High-Voltage Testing & Performance
Features, such as the ones showcased in subsection 3.3, could lead to localized field enhancement, which in turn could result in detrimental effects such as field emission and even breakdown [28, 30]. To assess the performance of an electrode, one cannot rely solely on electrostatic simulations, as the field around each feature is highly dependent on its geometry. Hence, to study such effects and to determine the overall performance of the electrode, we have performed a series of HV tests in a gaseous argon (GAr) environment. In those tests, we aimed to observe light associated with effects such as field emission or breakdown events.
@@ 4.1 High-Voltage Test Setup
A test setup, inspired by the one described in [86, 61], was built inside the Faraday cage. As illustrated in Figure 13, at the heart of the setup lies a square PVC box measuring $1.6\text{\,}\mathrm{m}$ in length and $20\text{\,}\mathrm{c}\mathrm{m}$ in depth. The box was elevated above the floor with the use of a dedicated four-legged frame with wheels. To maintain a GAr atmosphere and protect the electrode under test from environmental contaminants the box was instrumented with a 1.1 $1\text{\,}\mathrm{c}\mathrm{m}$ thick lid made of Makroclear® polycarbonate sheet [27], with strips of Armaflex® tape for sealing. The use of the polycarbonate lid transmits the light in the visible and in the near-infrared range [29, 54], allowing us to observe light, which could be associated with effects such as neutral bremsstrahlung electroluminescence and the avalanche scintillation of GAr in the near-infrared range [32, gAr_spectrum:2022]. The box was fitted with a GAr tube inlet, a HV connection to the electrode and a ground connection. The flow of GAr into the box was controlled via a flow meter connected to the bottles containing grade 5.0 argon. Additionally, an Infactory NC-7004-675 humidity and temperature sensor [48] was placed at the farthest corner from the GAr inlet, as these parameters can affect the breakdown voltage. Since the PVC box was not leak-tight, the pressure inside the box was assumed to be close to the atmospheric pressure, and thus not monitored.
@@
A $1.2\text{\,}\mathrm{m}$ in diameter and $5\text{\,}\mathrm{m}\mathrm{m}$ thick matt aluminum plate with a rounded edge was placed at the bottom of the box on top of a grounded SS plate, forming a ground plate in the setup. A number of PVC spacers was used to further support the aluminum plate on its sides, resulting in a total estimated elevation of $1\text{\,}\mathrm{c}\mathrm{m}$ of the plate above the box’s bottom. The maximum measured gap distance in between the electrode under test and the ground plate was $4.97\text{\,}\mathrm{c}\mathrm{m}$. This measured gap distance was used to define the so-called bulk field in subsequent analysis. However, it should be noted that throughout the HV tests the precise distance and alignment between the ground plate and the tested electrode were not controlled. The unevenness of the ground plate, defined as the vertical distance between its highest and lowest points, was measured to be $6.4\text{\,}\mathrm{m}\mathrm{m}$, amounting to about 15% of the gap distance.
The electrode mesh was mounted on the electrode frame in non-cleanroom environment conditions, with the use of the procedure described in subsection 3.2. However, as depicted in Figure 13 (top), for the purpose of the HV tests the stretching ring was not removed and the SS covers were not installed. The electrode was positioned in the setup with the side on which the mesh was mounted facing down towards the ground plate. To facilitate the HV connection, the electrode was connected to a connection in the box’s wall, which subsequently was connected to the toroid. For most of the tests, $20\text{\,}\mathrm{M\SIUnitSymbolOhm}$ resistors were added on the HV line leading from the test box to the toroid, limiting the current into the system and preventing potential electrode aging effects as a result of discharges [79]. In the final test configuration, the electrode frame was suspended from the box’s lid using PEEK [76] pillars wrapped in Kapton tape, as illustrated in Figure 13. Prior to that, several insulators with a varying geometry, produced from PVC, POM, and 3D-printed ASA plastic were tested as separators between the electrode frame and the ground plate. However, their use resulted in frequent breakdowns at the insulators, indicating that these insulators enhanced locally the field along the edge of the ground plate.